Consider the following data (in millions) from Trident Financial Inc., which has two main divisions, mortgage loans and consumer loans:

Mortgage Loans Consumer Loans
Average total assets $ 1,100 $ 21,250
Operating income $ 242 $ 2,550
Return on investment (ROI) 22.00 % 12.00 %

Required:

1. Based on ROI, which division is more successful?

  • Mortgage loans

  • Consumer loans

2. Trident uses residual income (RI) as a measure of the financial performance of its divisions. What is the RI for each division if the minimum desired rate of return is (a) 10%, (b) 15%, and (c) 20%? (Leave no cells blank. Enter "0" wherever required. Negative amounts should be indicated by a minus sign. Enter your answers in millions of dollars, rounded to nearest whole number.)

Solve:

Residual Income (RI) Mortgage Consumer Loans Loans (a) 10% (b) 15% (c) 20%
Answer #1

1. Mortgage Loans (as the return on investment is higher, it is more successful)

2. Residual income = Operating income - Required return

Mortgage loans Consumer loans
10%

= 242-(1100*10%)

= 132

= 2550-(21250*10%)

= 425

15%

= 242 - (1100*15%)

= 77

= 2550-(21250*15%)

= -638 (minus sign)

20%

= 242 - (1100*20%)

= 22

= 2550-(21250*20%)

= -1700 (minus sign)